Solve for the general solution for x: 4 cos^2 x-1=0
Found 2 solutions by lwsshak3, stanbon:
Answer by lwsshak3(11628)

Solve for the general solution for x:
4 cos^2 x-1=0
cos^2 x=1/4
cosx=1/2
x=π/3

Answer by stanbon(75887)

Solve for the general solution for x: 4 cos^2 x-1=0
-------
4cos^2(x) = 1
cos^2(x) = 1/4
cos(x) = 1/2
----
x = pi/3 or -pi/3
=========================
Cheers,
Stan H.
==============